Atterberg limits and natural water contents for five fine grained soils are listed in the following table:

Soil No.

LL (%)

PL (%)

Natural Water Content (%)

1

10

10

4

2

33

18

20

3

46

34

12

4

74

31

32

5

74

49

16

Classify the soils in the Unified Soil Classification System, estimate their optimum water contents and maximum dry densities, and indicate approximately how much each soil must be wetted or dried if it is to be efficiently compacted.
Answer #1

SOLUTLON N we have given fine grained soil. . 3011 Somple 1: Liquid limit (LL) = 10%. Plastic limit (PL) - 10% P.Inder (PL)-on plotting the given and obtained data in atter being graph or chand we get: A-line PE=0.73/11-20) Plasticity index 90 (PI)Samples: OMC = 0.910*10 -0.030 * 0 -0.875 = 0.29% MOD = -0.18 X 10 -0.027* 0+21182 = 19 since N. water cont. = 44 pamble in t